# VGPT-PINN: Viscosity-enhanced Generative Pre-Trained Physics Informed Neural Networks for parameterized nonlinear conservation laws

### v1

@@ 3.2 Reduced order model with a separable training process
  • At the first step, we freeze the output layer parameters $\{c_{i}(\mu)\}_{i=1}^{n}$ with $c_{i}(\mu)\equiv 1/n$ (or another initialization) and train the transform layer parameters $\{W_{\mu,\mu^{i}}\}_{i=1}^{n}$ and $\{b_{\mu,\mu^{i}}\}_{i=1}^{n}$ with a learning rate of $0.00005$. This step allows the network to identify and align shocks or discontinuities. The relatively larger updates to transform parameters allow the network focusing on capturing and accurately representing the locations and characteristics of these shocks, establishing a preliminary but reasonable approximation of the solution.
  • In the second step of the training process, we set different learning rates for the output layer $\{c_{i}(\mu)\}_{i=1}^{n}$ and the transform layer $\{W_{\mu,\mu^{i}}\}_{i=1}^{n}$ and $\{b_{\mu,\mu^{i}}\}_{i=1}^{n}$, at $0.005$ and $0.00005$ respectively.
@@
Algorithm 1 VGPT-PINN for parametric PDE: Online stage
@@
2:     Train the $n$-neuron VGPT-PINN at $\mu$ for all $\mu\in\Xi_{\rm train}$ with learning rate $lr=lr_{1}^{\text{in}}$ for the transform layer and $lr=lr_{1}^{\text{out}}$ for the output layer
3:end while
@@
5:     Train the $n$-neuron VGPT-PINN at $\mu$ for all $\mu\in\Xi_{\rm train}$ with learning rate $lr=lr_{2}^{\text{in}}$ for the transform layer and $lr=lr_{2}^{\text{out}}$ for the output layer
6:     Record the indicator $\Delta_{\mathsf{NN}}^{n}(\Theta(\mu))=\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}(\Theta(\mu))$ in Eq. (3.13)
@@
Output:  $n$-neuron VGPT-PINN $\mathsf{NN}^{n}(\mu)$
@@
By integrating these different learning rates, the VGPT-PINN framework effectively balances the need for rapid alignment of the discontinuities and the requirement for accurate representation of the solution. This approach enhances the network’s ability to capture both the global behavior and the local details, resulting in improved performance and higher accuracy in solving the problems.
Akin to the residual-based error estimation in traditional numerical solvers, our VGPT-PINN adopts the similar greedy offline process of the GPT-PINN [8]. Indeed, we utilize the terminal loss $\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}$ for each parameter as the error indicator $\Delta_{\mathsf{NN}}^{r}(\Theta(\mu))$ which allows us to incrementally expand the VGPT-PINN hidden layer, in a ground-up fashion, from zero to (a pre-determined) $N$ neurons or until certain stopping criteria are met (e.g. error indicator falling below a threshold). At each step, the parameter value that is most poorly approximated by the current meta-network is selected. A full PINN is then pre-trained to augment the hidden layer. In this fashion, the meta-network learns the system’s parametric dependencies one meta-neuron at a time. The offline training corresponds to the implementation of the greedy algorithm that is described in Algorithm 2.
Algorithm 2 VGPT-PINN for parametric PDE: Offline stage
Input: A random (or given) $\mu^{1}$, training set $\Xi_{\rm train}\subset\mathcal{D}$, and full PINN
1:Train a full PINN at $\mu^{1}$ to obtain $\Psi_{\mathsf{NN}}^{\mu^{1}}$. Precompute quantities necessary for $\nabla_{\Theta}(\mu)\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}$ at collocation nodes $\mathcal{C}_{o}^{r}$, $\mathcal{C}_{\partial}^{r}$, and $\mathcal{C}_{i}^{r}$. Set $n=2$
@@
3:     Train the ($n-1$)-neuron VGPT-PINN at $\mu$ for all $\mu\in\Xi_{\rm train}$ and record the indicator $\Delta_{\mathsf{NN}}^{r}(\Theta(\mu))$
@@
5:     Train a full PINN at $\mu^{n}$ to obtain $\Psi_{\mathsf{NN}}^{\mu^{n}}$. Precompute quantities necessary for $\nabla_{\Theta}\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}$ at collocation nodes $\mathcal{C}_{o}^{r}$, $\mathcal{C}_{\partial}^{r}$, and $\mathcal{C}_{i}^{r}$
6:     Update the VGPT-PINN by adding a neuron to the hidden VGPT-PINN layer to construct the $n$-neuron VGPT-PINN
@@
Output:  $N$-neuron VGPT-PINN with $N$ being the terminal index
@@ 4 Numerical results
In this section, we perform numerical results to demonstrate the capability of the VGPT-PINN in effectively approximating the Riemann problems of the Burgers’ and Euler equations parameterized by the initial conditions. Without relying on any prior knowledge or data, the method accurately captures, often using only one neuron, the complex dynamics and transitions of multiple shocks and rarefaction waves. The code for all these examples are published on GitHub at https://github.com/DuktigYajie/VGPT-PINN.
4.1 1D inviscid Burgers’ equation
For the 1D Burgers’ equation, our extensive tests include five cases as listed in Table 1. They are one single shock propagating by itself, two shocks becoming one, a smooth initial condition developing into a shock, one rarefaction wave, and finally a shock interacting with a rarefaction wave. They are named ${\rm B_{1S}}$, ${\rm B_{2S}}$, ${\rm B_{Sm}}$, ${\rm B_{R}}$ and ${\rm B_{RS}}$ respectively. The underlying PINNs have 5 hidden layers with 20 neurons each for all tests. They are trained with the Adam optimizer, using 30,000 epochs and an initial learning rate of 0.001. The space-time collocation set is a $99\times 199$ uniform grid excluding the boundaries. Moreover, 100 points each are placed to calculate the loss from the boundary and initial conditions.
The VGPT-PINN is trained using the Adam optimizer with a learning rate of 0.001 over 10,000 epochs.
Test name Initial condition $\phi(x,t=0)$ Boundary condition $g(\partial\Omega,t)$
@@
We pre-train a single PINN at $\mu=1$. The full-PINN solution and its comparison with the exact solution are shown in Figure 2 (a-c).
The one-neuron VGPT-PINN is tested on $\mu\in[1:0.025:2]$. We evaluate the L1 and L2 errors, and display them in Figure 2 (d,e). The variation of solutions, detailed errors and losses for two particular parameters unseen during training are shown in Figure 2 (f-m). Comparing with the single-PINN results, it is clear that our VGPT-PINN solutions achieve the same accuracy. Note that here the “smooth error” is calculated by excluding space-time points within a distance of $0.02$ from the discontinuity line.
Figure 2: Results of Case ${\rm B_{1S}}$. (a-c) Full PINN at $\mu=1.0$ and its comparison with the exact solution; (d-m) One-neuron VGPT-PINN error and error distribution; and its detailed results for $\mu=0.75$ (f-i), $\mu=1.75$ (j-m).
@@
We generate two independently-trained networks that are connected at the meeting time whose exact value $t=2/\mu_{2}$ is numerically predicted by the networks via (3.10). The first-stage is trained using a learning rate of 0.001 and 60000 epochs, while the second-stage has a learning rate 0.0001 and 40000 epochs. The VGPT-PINN uses the same temporal segmentation network structure as the PINN, dividing the time domain into segments with separate networks trained for each segment. This structure facilitates effective handling of time-dependent problems through division of the time domain into manageable segments, each with its own trained network. The VGPT-PINN matches PINN’s learning rates but requires less than one-tenth of the epochs. The results of both methods for two unseen parameters are shown in Figure 3. It is clear that the shocks and their interaction are accurately captured by our new method. Moreover, the VGPT-PINN errors match those of the PINN’s.
Figure 3: Results of Case ${\rm B_{2S}}$. (a-c) Full PINN results for $\mu=(0.5,1.6)$; (d-f) full PINN results for $(1.0,2.0)$; (g-j) one-neuron VGPT-PINN results for $\mu=(0.75,1.6)$; and (k-n) one-neuron VGPT-PINN results for $(0.85,1.84)$.
Case ${\rm B_{Sm}}$ – Indeed, even smooth and continuous initial conditions may lead to the formation of shocks. We consider the 1D inviscid Burgers’ equation with space domain $\Omega=[0,1]$, time domain $t\in[0,1]$ and parameter domain $\mathcal{D}=[0.5,1.5]\times[0.1,0.4]$, with a smooth initial value and periodic boundary condition. In this scenario, the exact solution cannot be expressed explicitly. To obtain a reference solution, we adopt the Godunov scheme [28] and solve the equation on a sufficiently fine mesh. The full-PINN solution and this reference solution are shown in Figure 4 (a-c). The VGPT-PINN solutions and errors under three unseen parameters are shown in Figure 4 (d-l).
Figure 4: Results of Case ${\rm B_{Sm}}$. Full PINN for $\mu=(1.0,0)$ (a-c), and the one-neuron VGPT-PINN results for $\mu=(0.75,0)$ (d-f), $\mu=(0.75,0.1)$ (g-i), $\mu=(1.25,0.4)$ (j-l).
@@
The full PINN solutions, errors, and the reduction of the loss function when $\mu=1$ are shown in Figure 5(a-c). Subsequently, we test the VGPT-PINN formed with this single neuron for $\mu\in[0.5:0.0125:1]$ and show the Loss, L1 error and L2 error in Figure 5 (d, e). The variation of solutions, errors and losses for two unseen parameters are shown in Panels (f-m) of this figure.
Figure 5: Results of Case ${\rm B_{R}}$. Full PINN for $\mu=1.0$ (a-c), one-neuron VGPT-PINN test error distribution (d,e), and more details for $\mu=0.7$(f-i) and $\mu=0.9$(j-m).
Figure 6: Results of Case ${\rm B_{RS}}$. Full PINN for $\mu=1.0$ (a-c). One-neuron VGPT-PINN results for $\mu=0.5$ (d-g), $\mu=0.9$ (h-k).
@@
The PINN solutions, errors, and the reduction of the loss function for equation at $\mu=1$ are shown in Figure 6 (a-c). The variation in VGPT-PINN’s solutions, errors and loss under two unseen parameters are shown in Figure 6 (d-k).
4.2 1D Euler equations
For the 1D Euler equations, we build the underlying PINNs featuring 6 hidden layers with 60 neurons each for two tests, the classical Sod and Lax problems. The VGPT-PINN with $N$-snapshots has single hidden layer with $N$ neurons, using the two step training in Algorithm 1 with Adam optimizer and an initial learning rate of 0.001.
@@
where $p_{1}\in[1.0,2.0].$ The collocation set consists of randomly selected 5,000 interior points. We place 100 points each for the initial and boundary conditions. Additionally, 100 RH collocation points are drawn from a uniform mesh of $100\times 200$ in the $X\times T$ space. The stopping criterion for the loss was set at $10^{-5}$ with a maximum of 30,000 epochs. In the underlying PINN, we set $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the loss function (3.1). The three components of the space-time PINN solution for $p_{1}=1.0$ are shown in Figure 7 (a-c). The PINN and exact solutions at final time $T=0.1$ for three different parameters are shown in Figure 7 (d-i).
Figure 7: PINN results for the Sod problem. (a-c) $x-t$ solution for $p_{1}=1.0$, (d-f) PINN solution at final time and its comparison with a reference solution, and (g-i) the decay of the overall loss for three different $p_{1}$ values (left to right).
In the VGPT-PINN process, we only need 3,000 interior points from a uniform mesh of $100\times 200$ in the $x\times t$ space and utilized 100 initial points, 100 boundary points, and 100 RH collocation points. We set $\lambda(x,t)\equiv 1,\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=0$ in the first stage’s loss function, and $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the second stage. We consider all the Sod problems with parameter domain $p_{1}=[1.0,2.0]$ and test the VGPT-PINN from 1 to 5 neurons. The approximate solution at $t=0.1$ will gradually approach the reference solution. Using just three snapshots, the approximate solution attains high accuracy, capturing the shock waves precisely. The VGPT-PINN solutions and loss are shown in Figure 8.
Figure 8: VGPT-PINN results for the Sod problem. Loss reduction and solution comparison for $\mu=1.3$ (left two columns) or $\mu=1.6$ (right two columns) for the VGPT-PINN generated by $1,2$ and $3$ neurons (from top to bottom).
@@
where $\rho_{1}\in[0.3,0.7].$ We set $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the loss function. The collocation set consists of randomly selected 30,000 interior points. Moreover, 1000 initial points, 1000 boundary points and 1000 RH collocations with Latin Hypercube Sampling method [31] in the $x\times t$ space, and the total loss was 0.01 after training for 20,000 epochs. The PINN solutions for Lax problem with $\rho=0.5$ are shown in Figure 9 (a-c). We also present the well-trained full-PINNs and losses for some parameters and compare them with the reference solution in Figure 9 (d-i).
@@
Figure 9: PINN results for the Lax problem. (a-c) $x-t$ solution for $\rho_{1}=0.5$, (d-f) PINN solution at final time and its comparison with a reference solution, and (g-i) the decay of the overall loss for three different $\rho_{1}$ values (left to right).
@@
In the VGPT-PINN process, we set $\lambda(x,t)\equiv 1,\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=0$ for the loss function in the first stage, and then $\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=10$ in the second stage. Using the unseen parameters $\rho=0.56$ and $\rho=0.618$, we show that as the number of neurons increases, the VGPT-PINN solution steadily approaches the reference solution in Figure 10. In the end, using just five neurons, the approximate solution attains high accuracy, capturing both the shock waves and shock contact precisely.
@@
Figure 10: VGPT-PINN results for the Lax problem: Loss reduction and solution comparison for $\mu=0.56$ (left two columns) or $\mu=0.618$ (right two columns) for the VGPT-PINN generated by $1,3$ and $5$ neurons (from top to bottom).
4.3 2D Euler equations
Finally, we consider a 2D example with strong shock [30]. A transonic flow with Mach number 0.728 passes through a stationary circular cylinder centered at $(1,1)$ with a parametric radius $r\in[0.2,0.3]$. We initialize the problem with a uniform flow of $(\rho,u,v,p)=(2,112,1.028,0,3.011)$. The computational domain is $(x,y,t)\in[0,1.5]\times[0,2]\times[0,0.4]$. The full PINN consists of 7 hidden layers, each with 90 neurons. We used a total of 300,000 collocation points obtained through the Latin hypercube sampling in the 3-dimensional space-time domain. Additionally, we randomly sample 15,000 boundary points from the cylinder’s surface and penalize the loss from the initial condition at 15,000 points. We set $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=0$ in both the PINN and VGPT-PINN losses. The PINN solutions at the final time $t=0.4$ are shown on the first row of Figure 11. The VGPT-PINN solutions for density $\rho$ and the overall loss with different neurons are present in the next two rows of Figure 11. And the absolute error for two unseen parameters $r=0.24$ and $r=0.26$ are shown in last two rows of Figure 11. Clearly, with 4 neurons, the density profile from VGPT-PINN matches that of the full PINN. Similar results were obtained for the pressure and velocities.
Figure 11: PINN and VGPT-PINN results for 2D Euler equations. (a-d): PINN solution at $t=0.4$ for the cylinder with $r=0.25$. Density profile (e-h) and the overall loss (i-l) for VGPT-PINN with different neurons. (m-v): the absolute VGPT-PINN error for two unseen parameters $r=0.24$ (i-m) and $r=0.26$.

### v2

@@ 3.2 Reduced order model with a separable training process
  • At the first step, we freeze the output layer parameters $\{c_{i}(\mu)\}_{i=1}^{n}$ with $c_{i}(\mu)\equiv 1/n$ (or another initialization) and train the transform layer parameters $\{W_{\mu,\mu^{i}}\}_{i=1}^{n}$ and $\{b_{\mu,\mu^{i}}\}_{i=1}^{n}$ with a learning rate of $10^{-5}$. This step allows the network to identify and align shocks or discontinuities. The relatively larger updates to transform parameters allow the network focusing on capturing and accurately representing the locations and characteristics of these shocks, establishing a preliminary but reasonable approximation of the solution.
  • In the second step of the training process, we set different learning rates for the output layer $\{c_{i}(\mu)\}_{i=1}^{n}$ and the transform layer $\{W_{\mu,\mu^{i}}\}_{i=1}^{n}$ and $\{b_{\mu,\mu^{i}}\}_{i=1}^{n}$, at $10^{-3}$ and $10^{-5}$ respectively.
@@
Algorithm 1 EGPT-PINN for parametric PDE: Online stage
@@
3:         Train the $n$-neuron EGPT-PINN at $\mu$ with learning rate $lr=lr_{1}^{\text{in}}$ for the transform layer and $lr=lr_{1}^{\text{out}}$ for the output layer
@@
5:              Train the $n$-neuron EGPT-PINN at $\mu$ with learning rate $lr=lr_{2}^{\text{in}}$ for the transform layer and $lr=lr_{2}^{\text{out}}$ for the output layer
6:              Record the indicator $\Delta_{\mathsf{NN}}^{n}(\Theta(\mu))=\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}(\Theta(\mu))$ in Eq. (3.13)
@@
8:     end while
9:end for
Output:  $n$-neuron EGPT-PINN $\mathsf{NN}^{n}(\mu)$
@@
By integrating these different learning rates, the EGPT-PINN framework effectively balances the need for rapid alignment of the discontinuities and the requirement for accurate representation of the solution. This approach enhances the network’s ability to capture both the global behavior and the local details, resulting in improved performance and higher accuracy in solving the problems.
Akin to the residual-based error estimation in traditional numerical solvers, our EGPT-PINN adopts the similar greedy offline process of the GPT-PINN [8]. Indeed, we utilize the terminal loss $\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}$ for each parameter as the error indicator $\Delta_{\mathsf{NN}}^{r}(\Theta(\mu))$ which allows us to incrementally expand the EGPT-PINN hidden layer, in a ground-up fashion, from zero to (a pre-determined) $N$ neurons or until certain stopping criteria are met (e.g. error indicator falling below a threshold). At each step, the parameter value that is most poorly approximated by the current meta-network is selected. A full PINN is then pre-trained to augment the hidden layer. In this fashion, the meta-network learns the system’s parametric dependencies one meta-neuron at a time. The offline training corresponds to the implementation of the greedy algorithm that is described in Algorithm 2.
Algorithm 2 EGPT-PINN for parametric PDE: Offline stage
Input: A random (or given) $\mu^{1}$, training set $\Xi_{\rm train}\subset\mathcal{D}$, and full EPINN
1:Train a full EPINN at $\mu^{1}$ to obtain $\Psi_{\mathsf{NN}}^{\mu^{1}}$. Set $n=2$
@@
3:     Train the ($n-1$)-neuron EGPT-PINN at $\mu$ for all $\mu\in\Xi_{\rm train}$ and record the indicator $\Delta_{\mathsf{NN}}^{r}(\Theta(\mu))$
@@
5:     Train a full PINN at $\mu^{n}$ to obtain $\Psi_{\mathsf{NN}}^{\mu^{n}}$.
6:     Update the EGPT-PINN by adding a neuron to the hidden EGPT-PINN layer to construct the $n$-neuron EGPT-PINN
@@
Output:  $N$-neuron EGPT-PINN with $N$ being the terminal index
@@ 4 Numerical results
In this section, we perform numerical results to demonstrate the capability of the EGPT-PINN in effectively approximating the Riemann problems of the Burgers’ and Euler equations parameterized by the initial conditions. Without relying on any prior knowledge or data, the method accurately captures, often using only one neuron, the complex dynamics and transitions of multiple shocks and rarefaction waves. The code for all these examples are published on GitHub at https://github.com/DuktigYajie/EGPT-PINN.
4.1 Inviscid Burgers’ equation
For the 1D Burgers’ equation, our extensive tests include five cases as listed in Table 1. They are one single shock propagating by itself, two shocks becoming one, a smooth initial condition developing into a shock, one rarefaction wave, and finally a shock interacting with a rarefaction wave. They are named ${\rm B_{1S}}$, ${\rm B_{2S}}$, ${\rm B_{Sm}}$, ${\rm B_{R}}$ and ${\rm B_{RS}}$ respectively. These five test cases represent basic formation and interaction examples of shocks and rarefaction waves, and are building blocks for more complicated scenarios.
Test name Initial condition $\phi(x,t=0)$ Boundary condition $\mathcal{G}(\partial\Omega,t)$
@@ 4.1 Inviscid Burgers’ equation
The underlying EPINNs have 5 hidden layers with 20 neurons each for all tests. They are trained with the Adam optimizer, using 25,000 epochs and an initial learning rate of 0.001. The space-time collocation set is a $99\times 199$ uniform grid excluding the boundaries. Moreover, 100 points each are placed to calculate the loss from the boundary and initial conditions. The EGPT-PINN is optimized using the Adam optimizer with a learning rate of 0.001 and trained over 1,000 epochs, which is less than one-tenth of the epochs required for the full EPINN training.
@@
We pre-train a single EPINN at $\mu=1$. The full-EPINN solution, its comparison with the exact solution, and the reduction of the loss function are presented in Figure 2 (a) attesting an accuracy of two significant digits.
The one-neuron EGPT-PINN is tested on $\mu\in[1:0.025:2]$. We evaluate the L1 and L2 errors across the parameter domain and present the results in Figure 3 (a). The variation of solutions, detailed errors and losses for two particular parameters unseen during training are shown in Figure 2 (b,c). Comparing with the full-EPINN results, it is clear that our EGPT-PINN solutions achieve the same accuracy. Note that here the “smooth error”, one order of magnitude lower than the whole domain error, is calculated by excluding space-time points within a distance of $0.02$ from the discontinuity line.
Figure 2: Results of Case ${\rm B_{1S}}$. (a) Full EPINN loss at $\mu=1.0$ and comparison with the exact solution; (b) One-neuron EGPT-PINN loss and comparison with the exact solution at $\mu=0.75$; (c) One-neuron EGPT-PINN loss and comparison with the exact solution at $\mu=1.75$.
Figure 3: Error test across the parameter domain for case ${\rm B_{1S}}$ (a) and ${\rm B_{R}}$ (b).
@@
We generate two independently-trained networks that are connected at the meeting time whose exact value $t=2/\mu_{2}$ is numerically predicted by the networks via (3.10). The first-stage is trained using a learning rate of 0.001 and 60,000 epochs, while the second-stage has a learning rate 0.0001 and 40,000 epochs. The EGPT-PINN uses the same temporal segmentation network structure as the EPINN, dividing the time domain into segments with separate networks trained for each segment. This structure facilitates effective handling of time-dependent problems through division of the time domain into manageable segments, each with its own trained network. The EGPT-PINN matches EPINN’s learning rates but requires less than one-tenth of the epochs. The results of EGPT-PINN for two unseen parameters are shown in Figure 4. It is clear that the shocks and their interaction are accurately captured by our new method. Moreover, the EGPT-PINN errors match those of the PINN’s with both achieving two digits of accuracy.
Figure 4: Results of Case ${\rm B_{2S}}$. (a-c) Two-neurons EGPT-PINN results for $\mu=(0.75,1.6)$; (d-f) Two-neurons EGPT-PINN results for $\mu=(0.85,1.84)$.
Case ${\rm B_{Sm}}$ – In fact, even smooth and continuous initial conditions can lead to shock formation. We consider the 1D inviscid Burgers’ equation with space domain $\Omega=[0,1]$, time domain $t\in[0,1]$ and parameter domain $\mathcal{D}=[0.5,1.5]\times[0.1,0.4]$, with a smooth initial value and periodic boundary condition. In this scenario, the exact solution cannot be explicitly expressed. To obtain a reference solution, we adopt the Godunov scheme [28] and solve the equation on a sufficiently fine mesh.
The full-EPINN solution and this reference solution are shown in Figure 5 (a,d). The EGPT-PINN solutions and errors under three unseen parameters are shown in Figure 5 (b,e) and (c,f). It is clear that the EGPT-PINN scheme captures the location of the shock very accurately, even when it is moving, by only one neuron. Moreover, we achieve better than 2-digit accuracy away from the shock.
Figure 5: Results of Case ${\rm B_{Sm}}$. (a, d) Full EPINN results for $\mu=(1.0,0)$ , and the one-neuron EGPT-PINN results for $\mu=(0.75,0.1)$ (b, e), $\mu=(1.25,0.4)$ (c, f).
@@
The full EPINN solutions, errors, and the reduction of the loss function when $\mu=1$ are shown in Figure 6(a). Subsequently, we test the EGPT-PINN formed with this single neuron for $\mu\in[0.5:0.0125:1]$ and show the L1 error and L2 error in Figure 3 (b). The variation of solutions, errors and losses for two random parameters are shown in Panels (b, c) of this figure. Once again, the single-neuron EGPT-PINN solutions achieve an accuracy comparable to that of the full PINN. The rarefaction of what is initially a shock is precisely captured as time evolves. We note that, had the entropy condition not been enforced, the accuracy of the solution would have deteriorated by multiple orders of magnitude. This underscores the robustness of the scheme due to its incorporation of the entropy condition.
Figure 6: Results of Case ${\rm B_{R}}$. Full EPINN results for $\mu=1.0$ (a), one-neuron EGPT-PINN results for two unseen parameters $\mu=0.7$ (b) and $\mu=0.9$ (c).
@@
The PINN solutions, errors, and the reduction of the loss function for equation at $\mu=1$ are shown in Figure 7 (a-c). The variation in EGPT-PINN’s solutions, errors and loss under two unseen parameters are shown in Figure 7 (d-k). These results demonstrate that the EGPT-PINN with one neuron is capable of accurately capturing two different kinds of waves that are simultaneously evolving.
Figure 7: Results of Case ${\rm B_{RS}}$. Full EPINN results for $\mu=1.0$ (a-c). One-neuron EGPT-PINN results for $\mu=0.5$ (d-g), $\mu=0.9$ (h-k).
4.2 Euler equations
For the 1D Euler equations, we build the underlying EPINNs featuring 6 hidden layers with 60 neurons each for two tests, the classical Sod and Lax problems. The EGPT-PINN with $N$-snapshots has single hidden layer with $N$ neurons, using the two step training in Algorithm 1 with Adam optimizer and an initial learning rate of 0.001.
@@ 4.2 Euler equations
where $p_{1}\in[1.0,2.0].$ The collocation set consists of randomly selected 5,000 interior points. We place 100 points each for the initial and boundary conditions. Additionally, 100 RH collocation points are drawn from a uniform mesh of $100\times 200$ in the $X\times T$ space. The stopping criterion for the loss was set at $10^{-5}$ with a maximum of 30,000 epochs. In the underlying EPINN, we set $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the loss function (3.1). The three components of the space-time EPINN solution for $p_{1}=1.0$ are shown in Figure 8 (a-c). The EPINN and exact solutions at final time $T=0.1$ for three different parameters are shown in Figure 8 (d-f).
Figure 8: EPINN results for the Sod problem. (a-c) $x-t$ solution for $p_{1}=1.0$, (d-f) EPINN solution at final time and its comparison with a reference solution.
In the EGPT-PINN process, we only need 3,000 interior points from a uniform mesh of $100\times 200$ in the $x\times t$ space and utilized 100 initial points, 100 boundary points, and 100 RH collocation points. We set $\lambda(x,t)\equiv 1,\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=0$ in the loss function of the first stage, and $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the second stage. We consider all Sod problems with parameter domain $[1.0,2.0]\ni p_{1}$ and test EGPT-PINNs with 1 to 3 neurons. As demonstrated by Figure 9, the approximate solution at $t=0.1$ gradually approaches the reference solution. Using just three neurons, the approximate solution attains high accuracy, capturing the shock waves precisely. The EGPT-PINN solutions and the loss are shown in Figure 9. It is clear by comparing with Figure 8 that, with only three neurons, the EGPT-PINN solutions achieve an accuracy that is indistinguishable from that of the full EPINNs.
Figure 9: EGPT-PINN results for the Sod problem. Loss reduction and solution comparison for $\mu=1.3$ (left two columns) or $\mu=1.6$ (right two columns) for the EGPT-PINN generated by $1,2$ and $3$ neurons (from top to bottom).
@@
where $\rho_{1}\in[0.3,0.7].$ We set $\varepsilon_{i}=\varepsilon_{b}=10$ and $\varepsilon_{r}=100$ in the loss function. The collocation set consists of randomly selected 30,000 interior points. Moreover, 1000 initial points, 1000 boundary points and 1000 RH collocations with Latin Hypercube Sampling method [31] in the $x\times t$ space, and the total loss was 0.01 after training for 20,000 epochs. The EPINN solutions for Lax problem with $\rho=0.5$ are shown in Figure 10 (a-c). We also present the well-trained full-EPINNs for some parameters and compare them with the reference solution in Figure 10 (d-f).
@@
Figure 10: EPINN results for the Lax problem. (a-c) $x-t$ solution for $\rho_{1}=0.5$, (d-f) EPINN solution at final time and its comparison with a reference solution.
@@
In the EGPT-PINN process, we set $\lambda(x,t)\equiv 1,\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=0$ for the loss function in the first stage, and then $\varepsilon_{i}=\varepsilon_{b}=1$ and $\varepsilon_{r}=10$ in the second stage. Using the unseen parameters $\rho=0.56$ and $\rho=0.618$, we show that as the number of neurons increases, the EGPT-PINN solution steadily approaches the reference solution in Figure 11. In the end, using just five neurons, the approximate solution attains high accuracy, capturing both the shock waves and shock contact precisely.
@@
Figure 11: EGPT-PINN results for the Lax problem: Loss reduction and solution comparison for $\mu=0.56$ (left two columns) or $\mu=0.618$ (right two columns) for the EGPT-PINN generated by $1,3$ and $5$ neurons (from top to bottom).
4.3 Inverse problem
To illustrate the robustness of EGPT-PINN, we task it to solve inverse problems. Toward that end, we consider the general formulation of an inverse problem: Given partial information about the solution, the objective is to infer unknown parameters or initial conditions of the underlying system. It has been widely demonstrated that EPINNs possess unique advantages in solving inverse problems. Traditional approaches often require multiple iterative steps, leading to significantly higher computational costs compared to forward problems. In contrast, neural network-based methods integrate the unknown parameters directly into the training process alongside network parameters, thereby achieving computational costs comparable to solving forward problems. The question we strive to answer in this paper is whether this advantage translates from the full EPINN to EGPT-PINN.
Taking the Sod shock tube problem as an example, suppose that we obtain measurements of density $\rho^{*}$, pressure $p^{*}$, and velocity $u^{*}$ at a specific location $(x^{*},t^{*})$ through experimental observations. The goal is to determine the corresponding initial conditions, such as the initial pressure. By incorporating the observed solution $\mathbf{U}^{*}=(\rho^{*},u^{*},p^{*})$ into the loss function and treating the initial conditions as trainable parameters, the network can simultaneously optimize the unknown parameters and fit the desired initial conditions during training, ultimately yielding an optimal solution. We achieve this by simply augmenting the forward loss function with a data-solution mismatch
$\mathcal{L}_{\mathrm{PINNInv}}^{\mathrm{EGPT}}=\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}+\varepsilon_{d}\left\|\mathbf{U}(x^{*},t^{*})-(\rho^{*},u^{*},p^{*})\right\|_{2},$ (4.3)
where $\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}$ the standard EGPT-PINN loss introduced in (3.14). In this setting, we consider the same configuration as in Section 4.2 for the Sod problem but assume that the initial pressure $p_{1}$ is unknown. Using EGPT-PINN, we simultaneously train both the parameterized solution $\Theta(\mu)$ and the unknown initial pressure $p_{1}$.
To test our inverse proglem solver, we select some $(x^{*},t^{*})$ and adopt the exact solution at the corresponding location as input data $\mathbf{U}^{*}$. We then train PINN (without RH-condition), EPINN and EGPT-PINN separately to predict the initial pressure $p_{1}$. The results are shown in Table 2 and Figure 12. The EGPT-PINN demonstrates remarkable effectiveness in accurately inferring the unknown initial condition, as evidenced by the significantly superior performance of both the predicted $p_{1}$ and the simultaneously obtained solution compared to those produced by the PINN and EPINN.
$(x^{*},t^{*})$ Reference $p_{1}$ PINN EPINN EGPT-PINN
(0.55,0.1) 1.3000 1.2995 1.2989 1.2998
(0.25,0.1) 1.3000 1.2993 1.2962 1.2995
(0.55,0.05) 1.6000 1.6003 1.5948 1.6005
(0.25,0.05) 1.6000 1.5990 1.5993 1.5998
Table 2: Four test cases for the inverse 1D Sod problem by the PINN, EPINN and EGPT-PINN.
Figure 12: The predicted solutions for the inverse problem, obtained using the PINN, EPINN and EGPT-PINN methods, are presented in the second row of Table 2.